OpenStudy (anonymous):

Pleasee helpp! STVU is an isosceles trapezoid. If SV=3x+1 and TU=x+21, Find the value of X?

OpenStudy (anonymous):



OpenStudy (anonymous):

x=10 ?

OpenStudy (anonymous):

If it's an isosceles trapezium, then the 2 non-parallel sides are equal aren't they? Which means that the diagonals are equal. Just equate the 2 expressions and solve for x.

OpenStudy (anonymous):

3x+1=x+21 -x -x 2x+1=21 -1 -1 2x/2=20/2 x=10 ?